OpenStudy (anonymous):

Solve the equation by factoring. x2 = x + 6

OpenStudy (nurali):

x2 = x + 6 x2-x-6=0 x2-3x+2x-6=0 x(x-3)+2(x-3)=0 (x-3)(x+2)=0 x-3=0 or x+2=0 x=3 or x=-2 x=3,-2